OpenStudy (anonymous):

Part A: Write the expression x2 - 18x + 45 as a product of two linear expressions. Show your work and justify each step. Part B: Rewrite x2 - 6x + 9 as a square of a linear expression. Part C: Do the expressions in parts A and B have a common factor? Justify your answer.

OpenStudy (anonymous):

oh where is said x2, I meant x^2

OpenStudy (fibonaccichick666):

When it says that, it just means to factor them like in this one \[x^2+2x+1=(x+1)(x+1)\] can you do it now?

OpenStudy (anonymous):

Um, I'm not really sure what that means.

OpenStudy (fibonaccichick666):

ok, have you heard of completing the square yet?

OpenStudy (anonymous):

I might have.

OpenStudy (anonymous):

Like, if you explain it, I think I know what it is.

OpenStudy (fibonaccichick666):

I double checked, yours are just factoring

OpenStudy (anonymous):

ok thank you!

OpenStudy (fibonaccichick666):

Check back after you watch it, and we can go over the rest of the problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay, I finished

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

alright, so do you have any ideas on how to start part A?

OpenStudy (fibonaccichick666):

@QuestionsWolf

OpenStudy (anonymous):

we find a number that adds to -18 and multiplies to 45

OpenStudy (fibonaccichick666):

good! so now, it's easiest to start with the factors of 45 can you tell me them?

OpenStudy (anonymous):

45 and 1, 5 and 9, 15 and 3,

OpenStudy (fibonaccichick666):

now, do any of those add to 18?

OpenStudy (anonymous):

15 and 3

OpenStudy (fibonaccichick666):

good, now the only issue is that we need a -18, any ideas on how we can get that?

OpenStudy (anonymous):

Have a number that is negative

OpenStudy (fibonaccichick666):

tell me the two numbers then? I don't quite understand your response

OpenStudy (anonymous):

Um, -15 and -3

OpenStudy (fibonaccichick666):

perfect! so then what does \(x^2-18x+45=?\)

OpenStudy (anonymous):

x^2 - 15 - 3 + 45

OpenStudy (fibonaccichick666):

uhm so go back and look at the factoring video or my original response. You want two things multiplying each other

OpenStudy (fibonaccichick666):

to = that polynomial

OpenStudy (anonymous):

Im not really sure what you are asking?

OpenStudy (fibonaccichick666):

remember this: When it says that, it just means to factor them like in this one x2+2x+1=(x+1)(x+1) can you do it now? So you want that format

OpenStudy (anonymous):

okay, let me see.

OpenStudy (anonymous):

x^2 - 15x - 3x + 45 (x^2 - 15x) + (-3x + 45) x(x - 15) + -3(x + -15) (x - 15)(x - 15) (x - 15)(x - 3)

OpenStudy (fibonaccichick666):

yup yup can't say I've ever seen it broken down that way, but that works

OpenStudy (anonymous):

Oh xD well thank you!

OpenStudy (fibonaccichick666):

so you just follow the same algorithm for part b

OpenStudy (fibonaccichick666):

really great idea for expanding it like that... I never would have thought of it

OpenStudy (anonymous):

Thank you, and thank you for your help! :)

OpenStudy (fibonaccichick666):

np if you have problems with the next part just message me